OpenStudy (anonymous):

Which of the following are solutions to the equation sinx cos x = 1/4? There can be more than one. PI/3 + nPI/2 PI/6 + nPI/2 PI/12 + nPI 5PI/12 + nPI

zepdrix (zepdrix):

\[\large \sin x \cos x = \frac{1}{4}\]Multiplying both sides by 4 will give us:\[\large 4 \sin x \cos x = 1\]

zepdrix (zepdrix):

This is another problem where we want to remember a trig identity, specifically, Sine Double Angle. \[\large 2\sin x \cos x = \sin 2x\]

zepdrix (zepdrix):

If we apply this to our problem, we have... \[\large 4 \sin x \cos x = 1\]\[\large 2 \cdot 2\sin x \cos x = 1\]\[\large 2 \cdot \sin 2x = 1\]

zepdrix (zepdrix):

\[\large \sin 2x = 1/2\]

zepdrix (zepdrix):

We'll need to solve for 2X before we can solve for X. Remember the reference angle for this one? \[\large \sin \theta = 1/2\] What values of theta give you a half? :D

OpenStudy (anonymous):

um sin2x?

OpenStudy (anonymous):

wait ...er 60?

zepdrix (zepdrix):

Ummmmm I think it's attttttttttt 30 degrees actually :o Since we're dealing with radians, we want the radian measure of 30 degrees. (pi/6) BUT, that is the reference angle for 2x! So now we want to solve for x. \[\large 2x=\frac{\pi}{6}\]Dividing both sides by 2 gives us,\[\large x=\frac{\pi}{12}\] If it's confusing, maybe go back over the stuff I posted :3 or ask questions.

OpenStudy (anonymous):

so what are the answers if its pi/12? 0.o

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

It's the option involving pi/12 :D

zepdrix (zepdrix):

im pretty sure there's only one of those XD

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

well you are wrong @zepdrix it was c and d .....so ..yah

zepdrix (zepdrix):

oh i didn't realize you could choose 2 of the 4... wish you woulda mentioned that :\ i woulda paid more attention.. soz